# A Hybrid TDMA/CSMA Protocol for Time-Sensitive Traffic in Robot Applications

### v1

@@ III-B2 Feature 2: Multi-Session Superframe with Flexible Time Slot Allocation
$T_{f}=T_{TDMA}+T_{CSMA}^{ctl}+T_{CSMA}^{gen}$
where $T_{TDMA}$, $T_{CSMA}^{ctl}$ and $T_{CSMA}^{gen}$ denote the durations of the time-slotted TDMA session, the CSMA-based control session, and a general-purpose CSMA session, respectively.
Figure 8: Superframe Structure of the IEEE 802.11-Compatible Hybrid TDMA/CSMA Protocol for Real-Time Robot Control with Deterministic and Flexible Access.
Session 1 (Time-slotted TDMA Session): The time-slotted TDMA session is exclusively reserved for the mission-critical traffic. Each participating node, referred to as an STA in IEEE 802.11 terminology, is assigned dedicated time slots within the superframes. The duration of the TDMA session is given by:
$T_{TDMA}=N_{TDMA}\times\tau_{TDMA}$
where $N_{TDMA}$ is the number of TDMA slots, and $\tau_{TDMA}$ is the duration of each slot, including the transmission time for critical packets and associated overhead. Each slot is allocated to a specific user through a time slot allocation described in the CSMA-based control session.
Session 2 (CSMA-based Control Session): The control session is dedicated to transmitting control and management messages for protocol coordination and network operations. Specifically, three messages are included: 1) IEEE 802.11 standard management messages, including association, authentication, and the integration of new STAs into the network; 2) PTP synchronization messages [22] (discussed in Feature 1); and 3) time-slot allocation directives. The TDMA server dynamically allocates the number ($N_{TDMA}$) and duration ($\tau_{TDMA}$) of slots for each client node based on real-time network conditions and robot application requirements. The TDMA server updates these assignments and clients configure schedules accordingly through the CSMA-based control session.
Session 3 (General-purpose CSMA Session): The general-purpose CSMA session is used for non-time-critical traffic, such as large-volume data traffic or event-driven traffic, which do not require stringent latency guarantees. Nodes access the medium via CSMA/CA, using IEEE 802.11 mechanisms for flexible, best-effort delivery.
@@ III-B3 Feature 3: Beacon’s NAV protection
To protect the TDMA session and the control session from interference, we use the NAV of the beacon frame to reserve the medium, i.e., prevent non-relevant STAs from initiating transmissions. The length of the beacon NAV is set as the exact duration of the TDMA session plus the CSMA-based control session:
$T_{NAV}=T_{TDMA}+T^{ctl}_{CSMA}$
@@ III-C Hybrid TDMA/CSMA Protocol Implementation
Our implementation is based on the Openwifi project [28], as shown in Fig. 9 This project divides IEEE 802.11 system functionality into hardware and software domains. The hardware domain, implemented on the FPGA, handles time-critical PHY and MAC functions, including modulation and CSMA/CA-based access control, with Verilog-implemented blocks for carrier sensing. The Analog Devices AD9361 RF frontend supports wireless communication with precise timing at a 61.44 MHz sampling rate. The Advanced eXtensible Interface (AXI) Bus connects software and hardware, providing a 600 MB/s channel for packet data and control signals. The software domain, hosted on the ARM processor running Linux, includes the Openwifi driver and mac80211 subsystem for protocol scheduling and network configuration. On the hardware side, we extend Openwifi project by adding three features into the MAC layer, namely 1) PTP synchronization, 2) packet timestamping, and 3) multi-queuing. On the software side, we develop a TDMA driver with a network manager API to realize packet transmission scheduling and TDMA slot allocations. In the following, we will discuss these novel implementations respectively.
@@ III-C1 PTP Synchronization implementation
In our system, the PTP synchronization messages (beacon and synchronization response message as discussed in Section III-B1), exchanged exclusively during the CSMA-based control session, carry 64-bit nanosecond-resolution timestamps. At reception of the $i$-th client, the hardware logic updates local clock offsets $o_{i}$ and propagation delay $d_{i}$ estimates, which are then applied to adjust transmission schedules $s_{i,j}^{l}$ for its subsequent TDMA slots. Specifically, $o_{i}$ and $d_{i}$ are returned to the TDMA driver for the calculation of $s_{i,j}^{l}$ as in Eq. 1.
@@ III-C2 Packet Timestamping and multi-queue mechanism
Furthermore, we use a multi-queue architecture to separate TDMA and CSMA traffic based on assigned port numbers at the robotic application layer. Specifically, if a packet is identified as TDMA, both the packet and its timestamp are placed into the TDMA queue and the corresponding timestamp queue. CSMA traffic is further classified into control management messages, placed in the CSMA-control queue for the control session, and general data packets, placed in the CSMA-data queue for the general-purpose session. All CSMA queues operate in a first-in-first-out manner. The multi-queue mechanism integrates directly with the protocol’s three-session superframe structure for deterministic delivery of mission-critical TDMA traffic and flexible non-critical CSMA traffic.
@@ III-C3 Network Manager API
The network manager is implemented as a user-space daemon to provide socket-based APIs for real‑time protocol configuration and control. Through these APIs, a robotic application can request adjustments to TDMA slot parameters such as the number of slots ($N_{TDMA}$), slot durations ($\tau_{TDMA}$), or reassigning slot ownership among active robots to better match changing communication loads and task priorities. When a configuration change is issued, the network manager constructs an updated allocation table, packages it into an IEEE 802.11 compliant management frame and transmits it during the CSMA based control session.
The middleware layer provides a publish–subscribe interface with the TDMA scheduler. Robotic applications register for slot based transmission by submitting a subscription message with explicit QoS parameters such as the maximum allowable deadline, the desired update frequency (e.g., 10 Hz steering updates), and the traffic priority level. The network manager matches these QoS requirements with the current network state to generate a TDMA slot map, which is then queued in the control session for broadcast to all nodes. In kernel space, the TDMA driver uses this slot map to configure high resolution timers to trigger packet dequeue operation at assigned slot boundaries.
@@ IV Performance Evaluation of the hybrid TDMA/CSMA Protocol
The comparison of non-critical traffic shown in the Fig.10(a) and 10(b) with their CSMA counterparts in Fig. 6 shows that the throughput of event-driven traffic and large-volume data traffic remains virtually unchanged. Specifically, the average throughput of non-critical flows fluctuates by less than 2% between the two protocols throughout the 1000-second test interval. This demonstrates that the implementation of TDMA-based protection for mission-critical traffic maintains compatibility with existing robotic data exchanges without degrading their operational performance.
The Fig. 10(c) shows the delivery status of mission-critical control packets. The definition of delivery status is the same as in Section II-B. With the hybrid TDMA/CSMA protocol, the missed deadline error rate is reduced by 93% compared to the conventional CSMA baseline. In a 1000-second experimental evaluation conducted under identical network conditions as the CSMA baseline, only 0.51% of packets experienced packet loss error or missed deadline error. Detailed error analysis shows that packet loss errors account for 98% of total failures, while missed deadline errors constitute only a negligible fraction of the total errors. Notably, the transmission errors in mission-critical control traffic in hybrid protocol are sparse and temporally isolated. This is different from the case in the CSMA where missed deadline errors tend to occur in bursts due to contention and unpredictable queuing delays.
@@
To evaluate the practical impact of the proposed hybrid protocol, we ran robot path-tracking simulations in the Robot Operating System (ROS) environment [33] using the PurePursuit trajectory following algorithm [34]. We design a simulation framework to compare the CSMA communication baseline with the proposed hybrid protocol in robot coordination scenarios.
@@
Figure 11: Robot trajectories (red) versus reference path (green) at 3 m/s, 6 m/s, and 9 m/s. Under CSMA, and the proposed Hybrid TDMA/CSMA Protocol (Hybrid)
@@
$\mathbf{r}_{k}=[x_{r}(k),y_{r}(k)]^{\top},\quad k=1,2,\dots,N,$
@@
$\mathbf{p}_{i}=[x(i),y(i)]^{\top},\quad i=1,2,\dots,M,$
@@
$\mathbf{r}_{k}=\arg\min_{\mathbf{r}\in\{\mathbf{r}_{1},\mathbf{r}_{2},\dots,\mathbf{r}_{N}\}}\|\mathbf{p}_{i}-\mathbf{r}\|.$
@@
${{e}_{d}}(i)=\|{{\mathbf{p}}_{i}}-{{\mathbf{r}}_{k}}\|\sin{{\theta}_{i}}.$
@@
To evaluate the performance of the proposed Hybrid TDMA/CSMA framework, we perform multiple path‑tracking simulations in a ROS environment that emulate real‑world navigation scenarios under varying speeds and network conditions. Fig. 12, 13, 14 show the absolute lateral tracking errors at different speeds for the CSMA baseline and the proposed Hybrid TDMA/CSMA framework. Each scenario was repeated 100 times to observe the distribution of lateral errors with variations from the ROS Gazebo physics engine’s stochastic effects.
@@
${{e}_{\text{traj,RMS}}}=\sqrt{\frac{1}{M}\sum\limits_{i=1}^{M}{\|{{\mathbf{p}}_{i}}-{{\widetilde{\mathbf{r}}}_{i}}{{\|}^{2}}}}.$
@@
Fig. 15 shows the RMS trajectory error of the robotic path-tracking experiment at different velocities. The results compare the proposed Hybrid TDMA/CSMA framework, the CSMA baseline, and an ideal PurePursuit algorithm baseline. The PurePursuit baseline represents an ideal path‑tracking controller without any communication delay and error. In this case, the system maintains minimal RMSE and closely follows the reference trajectory at all tested velocities. When command losses or missed deadlines are introduced by the communication emulation layer, the CSMA baseline shows a rapid RMSE increase as velocity rises, reaching about 25 m at 12 m/s. The Hybrid TDMA/CSMA framework maintains low RMSE, close to the ideal baseline, because it avoids burst sequences of consecutive missed deadlines and only incurs occasional errors caused by poor channel conditions. Fig. 15 also shows that the performance gain of the hybrid protocol is greater at higher velocities. In these cases, control commands require tighter timing to avoid cumulative trajectory deviations because a single error can trigger unrecoverable path‑tracking failures.
The path‑tracking evaluation provides a representative case for comparing the CSMA baseline and the Hybrid TDMA/CSMA framework, as similar requirements exist in many robotic applications. Systems such as surgical robots demand precise and timely control signals to execute fine movements, while drone swarms rely on synchronized command delivery to maintain coordinated flight. In both cases, delays or bursts of packet loss can cause severe and irreversible errors. By prioritizing time‑sensitive traffic and limiting burst losses, the Hybrid TDMA/CSMA framework enhances control accuracy and operational stability, making it broadly applicable to demanding real‑time robotic scenarios.
@@ V Conclusion and Future Work
Motivated by these observations, we present a hybrid TDMA/CSMA protocol to address command‑traffic congestion in time‑sensitive robotic applications. By combining TDMA’s deterministic slot scheduling with CSMA’s flexible access, the protocol guarantees collision‑free, low‑latency delivery of mission‑critical commands while supporting heterogeneous robotic traffic. The proposed protocol has three key features: (i) high‑precision PTP synchronization for sub‑microsecond slot alignment, (ii) a three‑session superframe with dynamic TDMA slot allocation based on real‑time network conditions and application demands, and (iii) beacon‑based NAV protection to reserve the medium for TDMA and control sessions against external interference. The proposed protocol also supports seamless integration with IEEE 802.11‑based wireless infrastructures for cost‑effective deployment and rapid scaling in heterogeneous robotic networks.
The proposed protocol implemented on a real-time Zynq‑7000 SoC‑based SDR platform and the experimental data collected from a real wireless environment were tested in ROS using the PurePursuit tracking algorithm. In high‑speed path tracking, it reduced RMSE by up to 90% and kept packet loss for latency‑sensitive control messages near zero compared to conventional CSMA baseline. It is also applicable to general real‑time robot control and various cooperative robotic systems that require precise and time‑critical command delivery.

### v2

@@ III-B2 Feature 2: Multi-Section Superframe with Flexible Time Slot Allocation
$T_{f}=T_{TDMA}+T_{CSMA}^{ctl}+T_{CSMA}^{gen}$ (5)
where $T_{TDMA}$, $T_{CSMA}^{ctl}$ and $T_{CSMA}^{gen}$ denote the durations of the time-slotted TDMA section, the CSMA-based control section, and a general-purpose CSMA section, respectively.
Figure 8: Superframe Structure of the IEEE 802.11-Compatible Hybrid TDMA/CSMA Protocol for Real-Time Robot Control
Superframe Section 1 (Time-slotted TDMA Section): The time-slotted TDMA section is exclusively reserved for the mission-critical traffic. Each participating node, referred to as an STA in IEEE 802.11 terminology, is assigned dedicated time slots within the superframes. The duration of the TDMA section is given by:
$T_{TDMA}=N_{TDMA}\times\tau_{TDMA}$ (6)
where $N_{TDMA}$ is the number of TDMA slots, and $\tau_{TDMA}$ is the duration of each slot, including the transmission time for critical packets and associated overhead. Each slot is allocated to a specific user through a time slot allocation described in the CSMA-based control section.
Superframe Section 2 (CSMA-based Control Section): The control section is dedicated to transmitting control and management messages for protocol coordination and network operations. Specifically, three messages are included: 1) IEEE 802.11 standard management messages, including association, authentication, and the integration of new STAs into the network; 2) PTP synchronization messages [21] (discussed in Feature 1); and 3) time-slot allocation directives. The TDMA server dynamically allocates the number ($N_{TDMA}$) and duration ($\tau_{TDMA}$) of slots for each client node based on real-time network conditions and robot application requirements. The TDMA server updates these assignments and clients configure schedules accordingly through the CSMA-based control section.
Superframe Section 3 (General-purpose CSMA Section): The general-purpose CSMA section is used for non-time-critical traffic, such as large-volume data traffic or event-driven traffic, which do not require stringent latency guarantees. Nodes access the medium via CSMA/CA, using IEEE 802.11 mechanisms for flexible, best-effort delivery.
@@ III-B3 Feature 3: Beacon’s NAV protection
To protect the TDMA section and the control section from interference, we use the NAV of the beacon frame to reserve the medium, i.e., prevent non-relevant STAs from initiating transmissions. The length of the beacon NAV is set as the exact duration of the TDMA section plus the CSMA-based control section:
$T_{NAV}=T_{TDMA}+T^{ctl}_{CSMA}$ (7)
@@ III-C Hybrid TDMA/CSMA Protocol Implementation
Our implementation is based on the Openwifi project [27], as shown in Fig. 9. This project divides IEEE 802.11 system functionality into hardware and software domains. The hardware domain, implemented on the FPGA, handles time-critical PHY and MAC functions, including modulation and CSMA/CA-based access control, with Verilog-implemented blocks for carrier sensing. The Analog Devices AD9361 RF frontend supports wireless communication with precise timing at a 61.44 MHz sampling rate. The Advanced eXtensible Interface (AXI) Bus connects software and hardware, providing a 600 MB/s channel for packet data and control signals. The software domain, hosted on the ARM processor running Linux, includes the Openwifi driver and mac80211 subsystem for protocol scheduling and network configuration. On the hardware side, we extend Openwifi project by adding three features into the MAC layer, namely 1) PTP synchronization, 2) packet timestamping, and 3) multi-queuing. On the software side, we develop a TDMA driver with a network manager API to realize packet transmission scheduling and TDMA slot allocations. In the following, we will discuss these novel implementations respectively.
@@ III-C1 PTP Synchronization implementation
In our system, the PTP synchronization messages (beacon and synchronization response message as discussed in Section III-B1), exchanged exclusively during the CSMA-based control section, carry 64-bit nanosecond-resolution timestamps. At reception of the $i$-th client, the hardware logic updates local clock offsets $o_{i}$ and propagation delay $d_{i}$ estimates, which are then applied to adjust transmission schedules $s_{i,j}^{l}$ for its subsequent TDMA slots. Specifically, $o_{i}$ and $d_{i}$ are returned to the TDMA driver for the calculation of $s_{i,j}^{l}$ as in Eq. 4.
@@ III-C2 Packet Timestamping and multi-queue mechanism
Furthermore, we use a multi-queue architecture to separate TDMA and CSMA traffic based on assigned port numbers at the robotic application layer. Specifically, if a packet is identified as TDMA, both the packet and its timestamp are placed into the TDMA queue and the corresponding timestamp queue. CSMA traffic is further classified into control management messages, placed in the CSMA-control queue for the control section, and general data packets, placed in the CSMA-data queue for the general-purpose section. All CSMA queues operate in a first-in-first-out manner. The multi-queue mechanism integrates directly with the protocol’s three-section superframe structure for deterministic delivery of mission-critical TDMA traffic and flexible non-critical CSMA traffic.
@@ III-C3 Network Manager API
The network manager is implemented as a user-space daemon to provide socket-based APIs for real‑time protocol configuration and control. Through these APIs, a robotic application can request adjustments to TDMA slot parameters such as the number of slots ($N_{TDMA}$), slot durations ($\tau_{TDMA}$), or reassigning slot ownership among active robots to better match changing communication loads and task priorities. When a configuration change is issued, the network manager constructs an updated allocation table, packages it into an IEEE 802.11 compliant management frame and transmits it during the CSMA based control section.
The middleware layer provides a publish–subscribe interface with the TDMA scheduler. Robotic applications register for slot based transmission by submitting a subscription message with explicit QoS parameters such as the maximum allowable deadline, the desired update frequency (e.g., 10 Hz steering updates), and the traffic priority level. The network manager matches these QoS requirements with the current network state to generate a TDMA slot map, which is then queued in the control section for broadcast to all nodes. In kernel space, the TDMA driver uses this slot map to configure high resolution timers to trigger packet dequeue operation at assigned slot boundaries.
@@ IV Performance Evaluation of the hybrid TDMA/CSMA Protocol
The comparison of non-critical traffic shown in Fig.10(a) and 10(b) with their CSMA counterparts in Fig. 6 shows that the throughput of event-driven traffic and large-volume data traffic remains virtually unchanged. Specifically, the average throughput of non-critical flows fluctuates by less than 2% between the two protocols throughout the 1000-second test interval. This demonstrates that the implementation of TDMA-based protection for mission-critical traffic maintains compatibility with existing robotic data exchanges without degrading their operational performance.
The Fig. 10(c) shows the delivery status of mission-critical control packets. The definition of delivery status is the same as in Section II-B. With the hybrid TDMA/CSMA protocol, the missed deadline error rate is reduced by 93% compared to the conventional CSMA baseline. In a 1000-second experimental evaluation conducted under identical network conditions as the CSMA baseline, only 0.51% of packets experienced packet loss error or missed deadline error. Detailed error analysis shows that packet loss errors account for 98% of total failures, while missed deadline errors constitute only a negligible fraction of the total errors. Notably, the transmission errors in mission-critical control traffic in hybrid protocol are sparse and temporally isolated. This is different from the case in CSMA where missed deadline errors tend to occur in bursts due to contention and unpredictable queuing delays.
@@
To evaluate the practical impact of the proposed hybrid protocol, we ran a robot path-tracking simulation in ROS environment [32] using the PurePursuit trajectory following algorithm [33]. We design a simulation framework to compare the CSMA communication baseline with the proposed hybrid protocol in robot coordination scenarios.
@@
Figure 11: Robot trajectories (red) versus reference path (green) at 3 m/s, 6 m/s, and 9 m/s under CSMA, and the proposed hybrid TDMA/CSMA protocol (Hybrid)
@@
$\mathbf{r}_{k}=[x_{r}(k),y_{r}(k)]^{\top},\quad k=1,2,\dots,N,$ (8)
@@
$\mathbf{p}_{i}=[x(i),y(i)]^{\top},\quad i=1,2,\dots,M,$ (9)
@@
$\mathbf{r}_{k}=\arg\min_{\mathbf{r}\in\{\mathbf{r}_{1},\mathbf{r}_{2},\dots,\mathbf{r}_{N}\}}\|\mathbf{p}_{i}-\mathbf{r}\|.$ (10)
@@
${{e}_{d}}(i)=\|{{\mathbf{p}}_{i}}-{{\mathbf{r}}_{k}}\|\sin{{\theta}_{i}}.$ (11)
@@
To evaluate the performance of the proposed hybrid TDMA/CSMA framework, we perform the path‑tracking simulation in a ROS environment that emulate real‑world navigation scenarios under varying speeds and network conditions. Fig. 12, 13, 14 show the absolute lateral tracking errors at different speeds for the CSMA baseline and the proposed hybrid TDMA/CSMA framework. Each scenario was repeated 100 times to observe the distribution of lateral errors with variations from the ROS Gazebo physics engine’s stochastic effects.
@@
${{e}_{\text{traj,RMS}}}=\sqrt{\frac{1}{M}\sum\limits_{i=1}^{M}{\|{{\mathbf{p}}_{i}}-{{\widetilde{\mathbf{r}}}_{i}}{{\|}^{2}}}}.$ (12)
@@
Fig. 15 shows the RMS trajectory error of the robotic path-tracking experiment at different velocities. The results compare the proposed hybrid TDMA/CSMA framework, the CSMA baseline, and an ideal PurePursuit algorithm baseline. The PurePursuit baseline represents an ideal path‑tracking controller without any communication delay and error. In this case, the system maintains minimal RMS error and closely follows the reference trajectory at all tested velocities. When command losses or missed deadlines are introduced by the communication emulation layer, the CSMA baseline shows a rapid RMS error increase as velocity rises, reaching about 25 m at 12 m/s. The hybrid TDMA/CSMA framework maintains low RMS error, close to the ideal baseline, because it avoids burst sequences of consecutive missed deadlines and only incurs occasional errors caused by poor channel conditions. Fig. 15 also shows that the performance gain of the hybrid protocol is greater at higher velocities. In these cases, control commands require tighter timing to avoid cumulative trajectory deviations because a single error can trigger unrecoverable path‑tracking failures.
The path‑tracking evaluation provides a representative case for comparing the CSMA baseline and the gybrid TDMA/CSMA framework, as similar requirements exist in many robotic applications. Systems such as surgical robots demand precise and timely control signals to execute fine movements, while drone swarms rely on synchronized command delivery to maintain coordinated flight. In both cases, delays or bursts of packet loss can cause severe and irreversible errors. By prioritizing time‑sensitive traffic and limiting burst losses, the hybrid TDMA/CSMA framework enhances control accuracy and operational stability, making it broadly applicable to demanding real‑time robotic scenarios.
@@ V Conclusion and Future Work
Motivated by these observations, we present a hybrid TDMA/CSMA protocol to address command‑traffic congestion in time‑sensitive robotic applications. By combining TDMA’s deterministic slot scheduling with CSMA’s flexible access, the protocol guarantees collision‑free, low‑latency delivery of mission‑critical commands while supporting heterogeneous robotic traffic. The proposed protocol has three key features: (i) high‑precision PTP synchronization for sub‑microsecond slot alignment, (ii) a three‑section superframe with dynamic TDMA slot allocation based on real‑time network conditions and application demands, and (iii) beacon‑based NAV protection to reserve the medium for TDMA and control sections against external interference. The proposed protocol also supports seamless integration with IEEE 802.11‑based wireless infrastructures for cost‑effective deployment and rapid scaling in heterogeneous robotic networks.
The proposed protocol implemented on a real-time Zynq‑7000 SoC‑based SDR platform and the experimental data collected from a real wireless environment were tested in ROS using the PurePursuit tracking algorithm. In high‑speed path tracking, it reduced RMS trajectory error by up to 90% and kept packet loss for latency‑sensitive control messages near zero compared to conventional CSMA baseline. It is also applicable to general real‑time robot control and various cooperative robotic systems that require precise and time‑critical command delivery.